OpenStudy (anonymous):

"the exponential function y=3(4)^x grows by a factor of 4 between x=1 and x=2 . what factor does it grow by between x=3 and x=4?"

OpenStudy (michele_laino):

if we call with f(x) your function, namely: \[\large f\left( x \right) = 3 \times {4^x}\] then we can write: \[\Large \begin{gathered} f\left( 3 \right) = 3 \times {4^3} = 3 \times 64 = ...? \hfill \\ f\left( 4 \right) = 3 \times {4^4} = 3 \times 256 = ...? \hfill \\ \end{gathered} \] now: what is: \[\Large \frac{{f\left( 4 \right)}}{{f\left( 3 \right)}} = ...?\]

OpenStudy (anonymous):

the answer would be 4

OpenStudy (michele_laino):

yes!

OpenStudy (anonymous):

thank you so much

OpenStudy (michele_laino):

thank you!